OpenStudy (amtran_bus):

The temp drops from 24c to 10c in 12 hours. What is the average temp change per hour.

OpenStudy (anonymous):

(24-10)/12 = 14/12 = 1.16 c per hour

OpenStudy (amtran_bus):

=slope